If sample mean = 54, sample standard deviation = 6, sample size = 25, confidence level = 90%, calculate the confidence interval for the estimate of the population mean.

Select one:

a. 51.95; 56.05

b. 54.8; 90.1

c. 50.6; 54.5

d. 6.0; 25.2


1
Expert's answer
2021-08-26T15:38:22-0400

"M = 54 \\

\sigma = 6 \\

n=25"

Two-sided confidence interval:

"CI = (M - \frac{Z_c \times \sigma}{\sqrt{n}}, M + \frac{Z_c \times \sigma}{\sqrt{n}})"

For 90 % confidence level

"Z_c=1.645 \\

CI = (54 - \frac{1.645 \times 6}{\sqrt{25}}, 54 + \frac{1.645 \times 6}{\sqrt{25}}) \\

=(54-1.974, 54+1.974) \\

= (52.026, 55.974)"

There is no right option.